OpenStudy (anonymous):

Exponential equation- Find exact solution and decimal approx. 49x+7^x+1-30=0

sam (.sam.):

Is it 49^x+7^(x+1)-30=0

OpenStudy (anonymous):

For decimal approx: I got 0.565

OpenStudy (anonymous):

solution: log(3)/log(7)

sam (.sam.):

\[49^x+7^{x+1}-30=0\] \[(7^{x})^{2}+7^{x+1}-30=0\] \[(7^{x})^{2}+7^{x}*7^{1}-30=0\] Let 7^x= y \[y^2+7y-30=0\] \[(y+10)(y-3)=0\] \[y=-10,3\] ------------------------------------- \[7^x=-10~~,~7^x=3\] \[\log _{7}(-10)=x~~,~~\log _{7}3=x\]

sam (.sam.):

\[\log _{7}3=x\] \[x=\frac{\log_{}3}{\log7}\]